OpenStudy (anonymous):

Solve using addition and multiplication principle.. 0.7n-15+n>2n-8-0.4n

OpenStudy (anonymous):

n > 70

OpenStudy (anonymous):

can you tell me how you got the answer. I am confused about this problem... thanks

OpenStudy (anonymous):

This probably will not help you to solve it. From Mathematica:\[\text{Reduce}\left[\frac{7}{10}n-15+n>2n-8-\frac{4}{10}n\right]=n>70 \]